# Implicit Dynamical Flow Fusion (IDFF) for Generative Modeling

### v1

@@ A.1.2 Training objective derivations
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,{\mathbf{x}}_{1},p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})}\big{\|}{\mathbf{v}}_{t}({\mathbf{x}})-\hat{{\mathbf{v}}}_{t}({\mathbf{x}})\big{\|}^{2}$ (12)
The CFM loss, as defined in equation 5, is equivalent to the FM loss up to a constant value, implying that $\mathcal{L}_{\text{FM}}(\theta)=\mathcal{L}_{\text{CFM}}(\theta)$. Based on this equivalence, we can express the training objective for approximating the vector field in equation 7, $\hat{f}({\mathbf{x}},t\mid{\mathbf{x}}_{1},{\mathbf{x}}_{0};\theta)$, as:
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{...}\big{\|}\hat{{\mathbf{v}}}_{t}({\mathbf{x}}\mid{\mathbf{x}}_{1};\theta)-{\mathbf{v}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\big{\|}^{2}$ (13)
To recover unbiased samples, we replace $\hat{u}(.)$ with equation 10 and $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1},{\mathbf{x}}_{0})$ with equation 9, resulting in:
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{...}\bigg{[}\bigg{\|}\frac{\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}}{1-t}-\frac{{\mathbf{x}}_{1}-{\mathbf{x}}}{1-t}\bigg{\|}^{2}\bigg{]}$ (14)
$=\mathbb{E}_{...}\big{[}\beta(t)^{2}\big{\|}\hat{x}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}_{1}\big{\|}^{2}\big{]}$ (15)
@@
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{...}\big{[}\beta(t)^{2}\big{\|}\hat{x}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}_{1}\big{\|}^{2}\big{]}$ (16)
To approximate $\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, we can employ a time-dependent score-based model, $\epsilon({\mathbf{x}},t;\theta)$, using a continuous loss function with a weighting schedule $\lambda(t)$. Since $\nabla\log p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$ approaches infinity as $t$ tends to 0 or 1, it is necessary to standardize the loss to maintain consistency over time. We set $\lambda(t)$ such that the target has zero mean and unit variance, predicting the noise added in sampling $x_{t}$ before multiplying by $\sigma_{t}=\sigma_{0}t(1-t)$. This leads to $\lambda(t)=\sigma_{t}=\sigma_{0}t(1-t)$, ensuring that the regression target for $\epsilon_{\theta}$ is distributed as $N(0,1)$. Notably, this loss function is independent of the $\gamma$ value, allowing us to use the same model with different $\gamma$ values during the sampling process without altering the minima. The loss is defined as:
$\mathcal{L}_{\text{IDFF}}^{2}(\theta)=\mathbb{E}_{...}\big{[}\lambda(t)^{2}\left\|\epsilon({\mathbf{x}},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\right\|^{2}\big{]}$ (17)
where $\lambda(\cdot)$ is a set of positive weights. With a proper choice of $\lambda(t)$, this loss is equivalent to the original DDPM loss up to a constant value, according to Theorem 1 of [Song et al., 2020b].
Combining the two components, the IDFF loss function is defined as:
@@
$\mathcal{L}_{\text{IDFF}}(\theta)=\mathbb{E}_{...}\big{[}\beta(t)^{2}\|\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}_{1}\|^{2}+\lambda(t)^{2}\left\|\epsilon({\mathbf{x}},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\right\|^{2}\big{]}$ (19)
A.1.3 Balancing the IDFF flow
@@
A.1.4 Implementation details
@@
A.2 Additional results for image generation experiment
@@
A.3 3D-attractors
In this experiment, we assess IDFF’s performance in generating trajectories of chaotic systems from scratch. We generate trajectories with $K=2000$ samples in each trajectory from 3D attractors, specifically the Lorenz and Rössler attractors, which are chaotic systems with nonlinear dynamics. The parameters for the Lorenz and Rössler models are set to ${\sigma=10,\rho=28,\beta=8/3}$ and ${a=.2,b=.2,c=5.7}$, respectively, to produce complex trajectories in 3D space. We then train the IDFF model based on these trajectories. To model each attractor, we use an MLP with two hidden layers of 128 dimensions and two separate heads for $\hat{{\mathbf{x}}}_{1}(.,t,k;\theta)$ and $\epsilon({\mathbf{x}},t,k;\theta)$. Additionally, we incorporate two separate embedding layers for embedding $t$ and $k$, which are directly concatenated with the first hidden layer of the MLP. The optimized IDFF successfully generates samples of these trajectories from scratch. The generated trajectories are shown in Figure 12. The quality of the results demonstrates that IDFF can successfully simulate the behaviors of highly nonlinear and nonstationary systems such as attractors.
@@ A.4 SST forecasting visualization
A.5 2D-simulated static data and time-series

### v2

@@ A.1.2 Training objective derivations
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{\|}\mathbf{v}_{t}(\mathbf{x}_{t})-\hat{\mathbf{v}}_{t}(\mathbf{x}_{t})\big{\|}^{2}$ (20)
The CFM loss, as defined in equation 5, is equivalent to the FM loss up to a constant value, implying that $\mathcal{L}_{\text{FM}}(\theta)=\mathcal{L}_{\text{CFM}}(\theta)$. Based on this equivalence, we can express the training objective for approximating the vector field in equation 7, $\hat{f}(\mathbf{x}_{t},t\mid\mathbf{x}_{1},\mathbf{x}_{0};\theta)$, as:
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{\|}\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1};\theta)-\mathbf{v}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})\big{\|}^{2}$ (21)
To recover unbiased samples, we replace $\hat{u}(.)$ with equation 10 and $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1},\mathbf{x}_{0})$ with equation 9, resulting in:
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\bigg{[}\bigg{\|}\frac{\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{t}}{1-t}-\frac{\mathbf{x}_{1}-\mathbf{x}_{t}}{1-t}\bigg{\|}^{2}\bigg{]}$ (22)
$=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{[}\beta(t)^{2}\big{\|}\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{1}\big{\|}^{2}\big{]}$ (23)
@@
$\mathcal{L}_{\text{IDFF}}^{1}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{[}\beta(t)^{2}\big{\|}\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{1}\big{\|}^{2}\big{]}$ (24)
To approximate $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$, we can employ a time-dependent score-based model, $\epsilon(\mathbf{x}_{t},t;\theta)$, using a continuous loss function with a weighting schedule $\lambda(t)$. Since $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ approaches infinity as $t$ tends to 0 or 1, it is necessary to standardize the loss to maintain consistency over time. We set $\lambda(t)$ such that the target has zero mean and unit variance, predicting the noise added in sampling $\mathbf{x}_{t}$ before multiplying by $\sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$. This leads to $\lambda(t)=\sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$, ensuring that the regression target for $\hat{\bm{\epsilon}}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$. Notably, this loss function is independent of the $\gamma$ value, allowing us to use the same model with different $\gamma$ values during the sampling process without altering the minima. The loss is defined as:
$\mathcal{L}_{\text{IDFF}}^{2}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{[}\lambda(t)^{2}\left\|\hat{\bm{\epsilon}}(\mathbf{x}_{t},t;\theta)-\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})\right\|^{2}\big{]}$ (25)
where $\lambda(\cdot)$ is a set of positive weights. With a proper choice of $\lambda(t)$, this loss is equivalent to the original DDPM loss up to a constant value, according to Theorem 1 of [Song et al., 2020b]. Combining the two components, the IDFF loss function is defined as:
@@
$\mathcal{L}_{\text{IDFF}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1},p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}\big{[}\beta(t)^{2}\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{1}\|^{2}+\lambda(t)^{2}\left\|\epsilon(\mathbf{x}_{t},t;\theta)-\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})\right\|^{2}\big{]}$ (27)
A.2 Balancing the IDFF flow
@@
A.3 IDFF training and sampling algorithms for time series
Algorithm 3 IDFF sampling algorithm for time-series data.
for $k$ in $\{1,...,K\}$ do
    if $k==1$ then
         ${\mathbf{x}}^{0}_{0}\sim\mathcal{N}(0,\textbf{I})$
    else
         ${\mathbf{x}}^{k}_{0}\sim\mathcal{N}({\mathbf{x}}^{k-1}_{1},\sigma_{0}\textbf{I})$
    end if
    for $t$ in $[0,1/\Delta t)$ do
         $\sigma_{t}\leftarrow\sigma_{0}\sqrt{t(1-t)}$
         ${\mathbf{w}}^{k}_{t}\leftarrow\sqrt{(1-\sigma^{2}_{t})}\frac{\hat{{\mathbf{x}}}_{1}({\mathbf{x}}_{t},t,k;\theta)-{\mathbf{x}}^{k}_{t}}{1-t}+\frac{2\gamma-1}{2}\sigma_{t}\hat{\epsilon}({\mathbf{x}}_{t},t,k;\theta)$
    end for
end for
return Samples $\{{\mathbf{x}}^{k}_{1}\}_{k=1}^{k=K}$
Algorithm 4 IDFF training algorithm for time-series data.
Input: dataset distribution $p_{1}({\mathbf{x}}_{1})$, initial distribution $p_{0}({\mathbf{x}}_{0})$, maximum sequence length $K$, bandwidth $\sigma_{0}$, a set of positive weights $\lambda(.)$ and $\beta(.)$, initialized networks $\hat{{\mathbf{x}}}_{1}(.;\theta)$ and $\hat{\bm{\epsilon}}(.;\theta)$,
while Training do
    if $K==1$ then
         $k=1$
         ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}_{1})$, ${\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}}_{0})$
    else
         $k\sim\{1,...,K\}$ with equal probability
         ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}^{k}_{1})$, ${\mathbf{x}}_{0}\sim p_{1}({\mathbf{x}}^{k-1}_{1})$
    end if
    $\pi\leftarrow\mathrm{OT}({\mathbf{x}}_{1},{\mathbf{x}}_{0})$
    $({\mathbf{x}}_{0},{\mathbf{x}}_{1})\sim\pi$
    $t\sim\mathcal{U}(0,1)$
    $\bm{\mu}_{t}\leftarrow t{\mathbf{x}}_{1}+(1-t){\mathbf{x}}_{0}$
    $\sigma_{t}\leftarrow\sigma_{0}\sqrt{t(1-t)}$
    ${\mathbf{x}}_{t}\sim\mathcal{N}(\bm{\mu_{t}},\sigma_{t}^{2}\textbf{I})$
    $\mathcal{L}_{\text{IDFF}}(\theta)\leftarrow\beta(t)^{2}\|\hat{{\mathbf{x}}}_{1}({\mathbf{x}}_{t},t,k;\theta)-{\mathbf{x}}_{1}\|^{2}+\lambda(t)^{2}\left\|\bm{\epsilon}({\mathbf{x}}_{t},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})\right\|^{2}$
end while
return $\{\hat{{\mathbf{x}}}_{1}(.;\theta),\hat{\epsilon}(.;\theta)\}$
A.4 Implementation details
@@
A.5 Additional results for image generation experiment
@@
A.6 3D-attractors
In this experiment, we assess IDFF’s performance in generating trajectories of chaotic systems from scratch. We generate trajectories with $K=2000$ samples in each trajectory from 3D attractors, specifically the Lorenz and Rössler attractors, which are chaotic systems with nonlinear dynamics. The parameters for the Lorenz and Rössler models are set to ${\sigma=10,\rho=28,\beta=8/3}$ and ${a=.2,b=.2,c=5.7}$, respectively, to produce complex trajectories in 3D space. We then train the IDFF model based on these trajectories. To model each attractor, we use an MLP with two hidden layers of 128 dimensions and two separate heads for $\hat{{\mathbf{x}}}_{1}(.,t,k;\theta)$ and $\epsilon({\mathbf{x}}_{t},t,k;\theta)$. Additionally, we incorporate two separate embedding layers for embedding $t$ and $k$, which are directly concatenated with the first hidden layer of the MLP. The optimized IDFF successfully generates samples of these trajectories from scratch. The generated trajectories are shown in Figure 12. The quality of the results demonstrates that IDFF can successfully simulate the behaviors of highly nonlinear and nonstationary systems such as attractors.
@@ A.7 SST forecasting visualization
A.8 2D-simulated static data and time-series